Calcium hydroxide and calcium sulfate are both used to prepare a saturated solution, S. Both
solids are only sparingly soluble, and the pH of S is found to be 11.6.
(i) Given that the Ksp for calcium hydroxide is 4.0 x 10-5
and the Ksp for calcium sulfate is
2.5 x 10-5
, determine the concentration of sulfate ions present in S.
1
Expert's answer
2020-06-08T15:50:43-0400

Solution.

"pH + pOH = 14"

"pOH = 14 - pH"

pOH = 2.4

"[OH^-] = 10^{-pOH}"

"[OH^-] = 3.98 \times 10^{-3} M"

"[Ca^{2+}][OH^-]^2 = Ksp(Ca(OH)2)"

"[Ca^{2+}] = (\frac{Ksp(CaOH)2}{[OH^-]^2}) = 2.53 M"

"[Ca^{2+}][SO4^{2-}] = Ksp(CaSO4)"

"[SO4^{2-}] = \frac{Ksp(CaSO4)}{[Ca^{2+}]} = 9.88 \times 10^{-6} M"

Answer:

"[SO4^{2-}] = 9.88 \times 10^{-6} M"